Maria says that sin A= disagree 12/18, cos A= 9/18, and tan A= 12/9. Do you agree or disagree? If you agree, explain why. If you disagree, explain what she did incorrectly and provide the correct solution.
1
Expert's answer
2020-05-21T16:35:53-0400

tanA=sinA/cosA

if I agree, so lets sinA =x

12/9 = x/(9/18)

so, x = 12/18, but we have, that sinA is not 12/18, it can't be true.

So I disagree with that question.